OpenStudy (anonymous):

Part 1: Change 4x – 3y = –15 to slope-intercept form. You must show all work. (2 points) Part 2: State the slope and y-intercept (2 points)

OpenStudy (mertsj):

Solve it for y. That is slope intercept form.

OpenStudy (anonymous):

y = mx + b this is the standard form of the slope..

OpenStudy (mertsj):

1. Subtract 4x from both sides 2. Divide both sides by -3

OpenStudy (anonymous):

This is what I have. 4x – 3y = –15 4x-4x-3y=-15-4x -3y=-15-4x -3/-3y=-15-4x/-3 y=-15-4x/-3

OpenStudy (mertsj):